I have a rectangle 36" x 55". If it were 42" wide, how long would it be? How do I find that? (sorry, dumb)
Answer by stanbon(75887)

I have a rectangle 36" x 55". If it were 42" wide, how long would it be? How do I find that?
-------
If you want the rectangles to be similar, use a proportion.
x/42 = 36/55
---
x = 42(36/55)
---
x = 27.5
---------
If you want to maintain equal area,
x = (36*55)/42 = 47.14
============================
Cheers
Stan H.
==========